. Suppose you have a 9V battery, a 2μF capacitor, and a 7.40μF capacitor. 

(a) Find the charge and energy stored if the capacitors are connected to the 

battery in series. (b) Do the same for a parallel connection.



1
Expert's answer
2021-12-06T15:58:27-0500

(a) Series capacitors


"\dfrac{1}{C_{TS}}=\dfrac{1}{C_1}+\dfrac{1}{C_2}=\dfrac{1}{2\times 10^{-6}F}+\dfrac{1}{7.4\times 10^{-6}F}"


"=\dfrac{4.7}{7.4\times 10^{-6}F}"


"Q_S=C_{TS}U=\dfrac{7.4\times 10^{-6}F}{4.7}\cdot9V"

"\approx14.17\times 10^{-6} C=14.17\mu C"

"E_S=\dfrac{C_{TS}U^2}{2}=\dfrac{7.4\times 10^{-6}\cdot(9V)^2}{2(4.7)}"

"\approx63.766\times 10^{-6} J=63.766\mu J"


(b) Parallel  capacitors


"C_{TP}=C_1+C_2=2\times 10^{-6}F+7.4\times 10^{-6}F"

"=9.4\times 10^{-6}F"

"Q_P=C_{TP}U=9.4\times 10^{-6}F\cdot9V"


"=84.6\times 10^{-6} C=84.6\mu C"

"E_P=\dfrac{C_{TP}U^2}{2}=\dfrac{9.4\times 10^{-6}\cdot(9V)^2}{2}"

"=0.3807\times 10^{-3} J=0.3807m J"